# Distribution Aligned Semantics Adaption for Lifelong Person Re-Identification

### v1

@@ 4.5 Limitations and Future Work
Our proposed LReID paradigm is based on the pre-trained Re-ID models, which capture robust and general human semantic knowledge after being trained on diverse and challenging pedestrian images at a large scale. Although we show that current pre-trained models have been capable of supplying useful human semantics for lifelong evolution, collecting more diverse and less biased data for better pre-training is expected to expand their capacity.
@@ 5 Conclusion
In this paper, we propose the Distribution Aligned Semantics Adaption (DASA) framework that is free of exemplars and old classifiers to effectively take advantage of the pre-trained model for LReID. We assume the Re-ID model pre-trained with large-scale pedestrian data can acquire robust and general human semantics as shared knowledge for lifelong application. To effectively adapt the pre-trained model to each application domain, we propose to perform distribution alignment by tuning BN and propose the lightweight Semantics Adaption (SA) module to aggregate learned semantics for better pedestrian representations. Extensive experiments show the superiority of the proposed framework for LReID in storage consumption and performance. It is expected to inspire more research to make full use of pre-trained models for effective and efficient LReID, making the landing application of Re-ID possible.

### v2

@@ 4.5 Limitations and Future Work
Our proposed LReID paradigm leverages the pre-trained models, which capture robust and general human semantic knowledge after being trained on diverse and challenging pedestrian images at a large scale. Although we show that current pre-trained models have been capable of supplying useful semantics for lifelong evolution, collecting more diverse and less biased data for better pre-training is expected to expand their capacity.
@@ 5 Conclusion
In this paper, we propose the Distribution Aligned Semantics Adaption (DASA) framework that is free of exemplars and old models to effectively take advantage of the pre-trained model for LReID. We assume the Re-ID model pre-trained with large-scale pedestrian data can acquire robust and general human semantics as shared knowledge for lifelong application. To effectively adapt the pre-trained model to each application domain, we propose to perform distribution alignment by tuning BN and propose the lightweight Semantics Adaption (SA) module to aggregate learned semantics for better pedestrian representations. Extensive experiments show the superiority of the proposed framework for LReID in storage consumption and performance. It is expected to inspire more research to make full use of pre-trained models for effective and efficient LReID, making the landing application of Re-ID possible.
[…]
Acknowledgements This work was supported in part by the NSFC Project (62176061), in part by the Shanghai Municipal Commission of Economy and Informatization (2021-GZL-RGZN-01033), in part by the Shanghai Research and Innovation Functional Program (17DZ2260900), and in part by the Shanghai Technology Development and Entrepreneurship Platform for Neuromorphic and AI SoC.